# An Objective Improvement Approach

### v2

@@ 5.4. Making Games Sharp and Improving
Naturally, not every game is improving, or even sharp. In this subsection, we first discuss how to almost surely make games sharp by adding sufficiently small random noise to the edge weights, and then discuss how to treat games that are not improving by assigning randomly chosen factors, with which the offsets of edges are weighted. Note that these are ‘global’ adjustments of the game that only need to be applied once, as it is the game that becomes sharp and improving, respectively.
@@
Before defining the gap of the game we recall that $\mathsf{val}_{\sigma}$ is the valuation of the joint strategy $\sigma$ (while $\mathsf{val}$ denotes the outcome of the optimisation problem). We use this valuation to argue that a small distortion of the edge weights cannot turn non-co-optimal strategies into co-optimal ones, and we define the gap of the game to reason about how small is small enough to retain non-co-optimality. The value of a joint strategy itself is useful for this, because it is much easier to access than the result of optimisation. As a valuation of a non-co-optimal joint strategy is no solution of a game, we will encounter negative offsets: they occur for exactly those edges, for which the associated inequation is not satisfied by $\mathsf{val}_{\sigma}$.
@@
Contraction and Gap of a Game
Given a game $\mathcal{G}$, we call its contraction $\lambda^{*}=\max\{\lambda_{e}\mid e\in E\}$. For a non-co-optimal joint strategy $\sigma$, we call the gap of $\sigma$ for $\mathcal{G}$ the value $\gamma_{\sigma}=-\min\{\mathsf{offset}(\mathsf{val}_{\sigma},e)\mid e\in E\}$. We call the gap of $\mathcal{G}$ the value $\gamma=\min\{\gamma_{\sigma}\mid\sigma$ is a non-co-optimal joint strategy}}\}}, i.e. the minimal such value.
Note that $\gamma_{\sigma}>0$ always holds for non-co-optimal strategies; for the minimal [^3] This skips over the case where all strategies are co-optimal, but that case is trivial to check and such games are trivial to solve, so that we ignore this case in this subsection. value $\gamma$, $\gamma>0$ thus always holds. Therefore, it is possible to obtain a better value by playing a different edge than the one selected by $\sigma$.
@@
Let $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ be a DPG with contraction $\lambda^{*}$ and gap $\gamma$, and let
$\mathcal{G}^{\prime}=(V_{\min},V_{\max},E,w^{\prime},\lambda)$ differ from $\mathcal{G}$ only in the edge weights such that, for all $e\in E$, $|w_{e}-w_{e}^{\prime}|<\frac{1-\lambda^{*}}{2}\gamma$ holds. Then a joint co-optimal strategy from $\mathcal{G}^{\prime}$ is also co-optimal for $\mathcal{G}$.
@@
We argue that the small weight disturbance, $|w_{e}-w_{e}^{\prime}|<\frac{1-\lambda^{*}}{2}\gamma$ for all $e\in E$, provides a small difference in the valuation that does not affect co-optimality. By using the definition of $\mathsf{val}_{\sigma}$ and the triangle inequality we can estimate the difference between the valuations. Precisely, for all joint strategies $\sigma$, we have for $\mathsf{val}$ on $\mathcal{G}$, and $\mathsf{val}^{\prime}$ on $\mathcal{G}^{\prime}$, that $|\mathsf{val}_{\sigma}(v)-\mathsf{val}_{\sigma}^{\prime}(v)|<\frac{1}{1-\lambda^{*}}\frac{1-\lambda^{*}}{2}\gamma=\frac{\gamma}{2}$. In fact, $\sigma$ defines a play $\rho=e_{0}e_{1}e_{2}\ldots$, and we have $\mathsf{val}_{\sigma}(v)=\mathsf{out}(\rho)=\sum_{i=0}^{\infty}w_{e_{i}}\prod_{j=0}^{i-1}\lambda_{e_{j}}$ and $\mathsf{val}_{\sigma}^{\prime}(v)=\sum_{i=0}^{\infty}w_{e_{i}}^{\prime}\prod_{j=0}^{i-1}\lambda_{e_{j}}$. This provides
@@
$\begin{array}[]{cl}|\mathsf{offset}^{\prime}(\mathsf{val}_{\sigma}^{\prime},e)-\mathsf{offset}(\mathsf{val}_{\sigma},e)|&\leq|\mathsf{val}_{\sigma}^{\prime}(v)-\mathsf{val}_{\sigma}(v)|+|w_{e}-w_{e}^{\prime}|+\lambda_{e}|\mathsf{val}_{\sigma}(v^{\prime})-\mathsf{val}_{\sigma}^{\prime}(v^{\prime})|\\ &<\frac{\gamma}{2}+\frac{1-\lambda^{*}}{2}\gamma+\lambda_{e}\frac{\gamma}{2}\leq\gamma\\ \end{array}$
@@
Note that these distributions are continuous. They could be replaced by any other distribution over these intervals that has 00 weight for all individual points.
@@
In this proof, we show a stronger property: every two different bases almost surely define different valuations (regardless of whether or not these valuations are solutions). There are only finitely many bases, and it suffices to compare two arbitrary but fixed ones, say $b_{1}$ and $b_{2}$.
What we show is that it happens with probability 00 that $b_{1}$ and $b_{2}$ are bases and define the same solution. If $b_{1}$ or $b_{2}$ is not a basis, we are done, too.
As $b_{1}$ and $b_{2}$ are different, there will be one edge $e=(v,v^{\prime})$ that occurs in $b_{1}$, but not in $b_{2}$. As all weight disturbances are drawn independently, we assume without loss of generality that the weight disturbance to this edge is drawn last.
Now, the valuation $\mathsf{val}_{1}$ defined by $b_{1}$ does not depend on this final draw. For $\mathsf{val}_{1}$, there is a value $w_{e}^{\prime}=\mathsf{val}_{1}(v)-\lambda_{e}\mathsf{val}_{1}(v^{\prime})$ that defines the weight $w_{e}^{\prime}$ that $e$ would need to have, in order to make the inequation sharp for $e$.
For the valuation $\mathsf{val}_{1}$ defined by $b_{1}$ to be equal to the valuation $\mathsf{val}_{2}$ defined by $b_{2}$, the weight for the edge $e$ (after adding the drawn distortion) needs to be exactly $w_{e}^{\prime}$. (This is a necessary condition, but not necessarily a sufficient condition.) There is at most one value for the disturbance that would provide for this, and this disturbance for the weight of $e$ is sampled with a likelihood of 00.
Putting these two results together, we get:
@@
Given a pair of DPGs $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ with contraction $\lambda^{*}$ and gap $\gamma$, and $\mathcal{G}^{\prime}=(V_{\min},V_{\max},E,w^{\prime},\lambda)$ obtained from $\mathcal{G}$ by adding independently uniformly at random drawn values from an interval $(-\varepsilon,\varepsilon)$ to every edge weight, for some $\varepsilon\leq\frac{1-\lambda^{*}}{2}\gamma$, then, a joint co-optimal strategy from $\mathcal{G}^{\prime}$ is also co-optimal for $\mathcal{G}$, and $\mathcal{G}^{\prime}$ is almost surely sharp.
Note that we can estimate the gap cheaply when all coefficients in $\mathcal{G}$ are rational. The gap is defined as the minimum of $\gamma_{\sigma}$ over non-co-optimal joint strategies $\sigma$, and we start by fixing such a joint strategy.
For a given $v\in V$, $\sigma$ defines a run $\rho=e_{0}e_{1}e_{2}\ldots$ in the form of a “lasso path”, which consists of a (possibly empty) initial path $e_{0},\ldots,e_{k}$, followed by an infinitely often repeated cycle $e_{0}^{\prime},\ldots,e_{\ell}^{\prime}$, where the only vertex that occurs twice on the path $e_{0},\ldots,e_{k},e_{0}^{\prime},\ldots,e_{\ell}^{\prime}$ is the vertex reached before $e_{0}^{\prime}$ and after $e_{\ell}^{\prime}$, while all other vertices occur only once.
In this run, an edge $e_{i}$ occurs only once and contributes $\prod_{j=0}^{i-1}\lambda_{e_{j}}w_{e_{i}}$ to the value of this run, while an edge $e_{i}^{\prime}$ occurs infinitely often and contributes the value $\frac{(\prod_{j=0}^{k}\lambda_{e_{j}})\cdot(\prod_{j=0}^{i-1}\lambda_{e_{j}^{\prime}})\cdot w_{e_{i}^{\prime}}}{1-\prod_{j=0}^{\ell}\lambda_{e_{j}^{\prime}}}$ to the value of the run. Now all we need to do is to estimate a common denominator.
To do this, let $\mathsf{denom}(r)$ be the denominator of a rational number $r$. It is easy to see that
@@
is a common denominator of all expressions of the first type ($\prod_{j=0}^{i-1}\lambda_{e_{j}}w_{e_{i}}$ and of the nominator from the second ($(\prod_{j=0}^{k}\lambda_{e_{j}})\cdot(\prod_{j=0}^{i-1}\lambda_{e_{j}^{\prime}})\cdot w_{e_{i}^{\prime}}$).
This leaves the contribution of the denominator of the fraction from the second form, which is the nominator of the terms $1-\prod_{j=0}^{\ell}\lambda_{e_{j}^{\prime}}$. But as the term is between 00 and $1$, its value is strictly smaller than the value of its denominator; we thus have
$\mathsf{nom}(1-\prod_{j=0}^{\ell}\lambda_{e_{j}^{\prime}})\leq\prod_{j}^{l}\mathsf{denom}(\lambda_{e_{j}^{\prime}}).$
Considering that the cycles do not share states (or edges), we in particular have that smallest common multiple $\mathsf{scp}$ of these nominators is smaller than the product of all relevant edge denominators; we in particular have
$\mathsf{scp}\leq\prod\limits_{v\in V}\mathsf{denom}(\lambda_{(v,\sigma(v))})\ ,$
such that
$\mathsf{common}=\mathsf{common^{\prime}}\cdot\mathsf{scp}<\prod\limits_{v\in V}\mathsf{denom}(\lambda_{(v,\sigma(v))})^{2}\cdot\mathsf{denom}(w_{(v,\sigma(v))})$
is a common denominator for all occurring values.
Looking at an edge $e$ that defines $\gamma_{\sigma}$, then $\gamma_{\sigma}$ can be written with the denominator
$\mathsf{common}\cdot\mathsf{denom}(\lambda_{e}w_{e})\ .$
Obviously, the nominator is at least $1$.
Estimating $\gamma$ can thus be done by using the highest possible denominators available in $\mathcal{G}$. The representation of the resulting estimate $\gamma$ is polynomial in the size of $\mathcal{G}$.
@@
Let $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ be an improving DPG for a given set of positive numbers $\{\alpha_{e}>0\mid e\in E\}$, $\sigma$ a strategy for both players, $\mathsf{val}$ an optimal solution returned at Line 5 of Algorithm 3 for the adjusted function $f_{\sigma}^{\prime}$, and let there be no local improvements of $\sigma$ for $\mathsf{val}$. Then there is neighbouring valuation $\mathsf{val}^{\prime\prime}$ to $\mathsf{val}$ such that there is a better strategy $\sigma^{\prime}$ that satisfies $f_{\sigma^{\prime}}(\mathsf{val}^{\prime\prime})<f_{\sigma}(\mathsf{val})$.
Such a strategy $\sigma^{\prime}$ is better than $\sigma$, and it can be selected in a way that $(v,\sigma^{\prime}(v))\in E_{\mathsf{val}}^{\sigma}$ holds for all $v\in V$ for a given set $E_{\mathsf{val}}^{\sigma}$.
The proof from Theorem 5 can be repeated verbatim with the new offset definition.
@@
If each offset factor $\alpha_{e}$ is selected independently uniformly at random from a bounded interval of positive numbers444or from any other distribution over positive numbers that has 00 weight for all individual points, then a sharp DPG $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ is almost surely improving for a sampled set of positive numbers $\{\alpha_{e}>0\mid e\in E\}$.
@@
We show the claim for two arbitrary but fixed valuations $\mathsf{val}_{1}$ and $\mathsf{val}_{2}$ defined by two different bases, $b_{1}$ and $b_{2}$, respectively, that satisfy the inequations in $H$, and an arbitrary but fixed joint strategy $\sigma$. As there are finitely many bases and finitely many joint strategies, satisfying the requirement almost surely for them entails that the requirement is satisfied almost surely for the game.
As $\mathcal{G}$ is sharp, we have $\mathsf{val}_{1}\neq\mathsf{val}_{2}$.
@@
We therefore have that $\mathsf{offset}(\mathsf{val}_{1},(v,\sigma(v)))\neq\mathsf{offset}(\mathsf{val}_{2},(v,\sigma(v)))$ holds for some $v\in V$. As the $\alpha_{e}$ are drawn independently, we can assume w.l.o.g. that $\alpha_{(v,\sigma(v))}$ is drawn last. There is at most one value $\alpha_{(v,\sigma(v))}^{\prime}$ for which the following condition is not satisfied:
@@
It therefore almost surely holds for all strategies that all basis-induced valuations have a pairwise distinct image by the objective function associated to the strategy. This immediately implies that the game is improving.
First, it is sharp by assumption (the factors $\alpha_{e}$ do not affect sharpness, as they only affect the objective functions $f_{\sigma}$). Second, assume that a basis defines a solution $\mathsf{val}$ that is not optimal w.r.t. the adjusted $f_{\sigma}$, then the simplex algorithm will pick a neighbouring basis that defines a solution $\mathsf{val}^{\prime}$ with $f_{\sigma}(\mathsf{val}^{\prime})\leq f_{\sigma}(\mathsf{val})$. Finally, as we have just shown that $f_{\sigma}(\mathsf{val}^{\prime})\neq f_{\sigma}(\mathsf{val})$ holds almost surely and consider one of the games where this holds, this implies $f_{\sigma}(\mathsf{val}^{\prime})<f_{\sigma}(\mathsf{val})$, which is the definition of improving for sharp games.

### v3

@@ 5.4. Making Games Sharp and Improving
Naturally, not every game is sharp or improving. In this subsection, we first discuss how to almost surely make games sharp by adding sufficiently small random noise to the edge weights, and then discuss how to treat games that are not improving by assigning randomly chosen factors, with which the offsets of edges are weighted. Note that these are ‘global’ adjustments of the game that only need to be applied once, as it is the game that becomes sharp and improving, respectively.
@@
Before defining the gap of the game, we recall that $\mathsf{val}_{\sigma}$ is the value of the joint strategy $\sigma$ (whereas $\mathsf{val}$ denotes the outcome of the optimisation problem). We use this valuation to argue that a small distortion of the edge weights cannot turn non-co-optimal strategies into co-optimal ones, and we define the gap of the game to reason about how small is small enough to retain non-co-optimality. The value of a joint strategy itself is useful for this because it is much easier to access than the result of optimisation. As a valuation of a non-co-optimal joint strategy is not a solution of a game, we will encounter negative offsets: they occur for exactly those edges, for which the associated inequation is not satisfied by $\mathsf{val}_{\sigma}$.
@@
(Contraction and Gap of a Game) Given a game $\mathcal{G}$, we call its contraction $\lambda^{*}=\max\{\lambda_{e}\mid e\in E\}$. For a non-co-optimal joint strategy $\sigma$, we call the gap of $\sigma$ for $\mathcal{G}$ the value $\gamma_{\sigma}=\min\{-\mathsf{offset}(\mathsf{val}_{\sigma},e)\mid e\in E,\mathsf{offset}(\mathsf{val}_{\sigma},e)<0\}$. We call the gap of $\mathcal{G}$ the value $\gamma=\min\{\gamma_{\sigma}\mid\sigma$ is a non-co-optimal joint strategy}}\}}, i.e. the minimal such value.
Note that $\gamma_{\sigma}>0$ always holds for non-co-optimal strategies; for the minimal [^6] This skips over the case where all strategies are co-optimal, but that case is trivial to check and such games are trivial to solve, so that we ignore this case in this subsection. value $\gamma$, $\gamma>0$ thus always holds.
In a strategy improvement approach, where there is an edge $e$ with $-\mathsf{offset}(\mathsf{val}_{\sigma},e)>0$, this would constitute a profitable switch, allowing the player to improve their return. While we do not use it this way, we argue that the gap $\gamma$ allows us to infer that, when we change all edge weights by a sufficiently small factor (which we can derive from $\gamma_{\sigma}$, or $\gamma$, and the contraction of the game), a non-co-optimal strategy $\sigma$ is still non-co-optimal after adjusting the weights this way.
@@
Let $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ be a DPG with contraction $\lambda^{*}$ and gap $\gamma$, and let $\mathcal{G}^{\prime}=(V_{\min},V_{\max},E,w^{\prime},\lambda)$ differ from $\mathcal{G}$ only in the edge weights such that, for all $e\in E$, $|w_{e}-w_{e}^{\prime}|<\frac{1-\lambda^{*}}{2}\gamma$ holds. Then any joint co-optimal strategy from $\mathcal{G}^{\prime}$ is also co-optimal for $\mathcal{G}$.
@@
We argue that the small weight disturbance, $|w_{e}-w_{e}^{\prime}|<\frac{1-\lambda^{*}}{2}\gamma$ for all $e\in E$, provides a small difference in the value that does not affect co-optimality. By using the definition of $\mathsf{val}_{\sigma}$ and the triangle inequality we can estimate the difference between the values. Precisely, for all joint strategies $\sigma$, we have for $\mathsf{val}_{\sigma}$ on $\mathcal{G}$, and $\mathsf{val}_{\sigma}^{\prime}$ on $\mathcal{G}^{\prime}$, that $|\mathsf{val}_{\sigma}(v)-\mathsf{val}_{\sigma}^{\prime}(v)|<\frac{1}{1-\lambda^{*}}\frac{1-\lambda^{*}}{2}\gamma=\frac{\gamma}{2}$. Indeed, $\sigma$ defines a play $\rho=e_{0}e_{1}e_{2}\ldots$, and we have $\mathsf{val}_{\sigma}(v)=\mathsf{out}(\rho)=\sum_{i=0}^{\infty}w_{e_{i}}\prod_{j=0}^{i-1}\lambda_{e_{j}}$ and $\mathsf{val}_{\sigma}^{\prime}(v)=\sum_{i=0}^{\infty}w_{e_{i}}^{\prime}\prod_{j=0}^{i-1}\lambda_{e_{j}}$. This provides
@@
$\begin{array}[]{cl}|\mathsf{offset}^{\prime}(\mathsf{val}_{\sigma}^{\prime},e)-\mathsf{offset}(\mathsf{val}_{\sigma},e)|\!\!\!\!&\leq|\mathsf{val}_{\sigma}^{\prime}(v)-\mathsf{val}_{\sigma}(v)|+|w_{e}-w_{e}^{\prime}|+\lambda_{e}|\mathsf{val}_{\sigma}(v^{\prime})-\mathsf{val}_{\sigma}^{\prime}(v^{\prime})|\\ &<\frac{\gamma}{2}+\frac{1-\lambda^{*}}{2}\gamma+\lambda_{e}\frac{\gamma}{2}\leq\gamma.\\ \end{array}$
@@
Note that these distributions are continuous. They could be replaced by any other distribution over these intervals that has weight 00 for all individual points.
@@
In this proof, we show a stronger property: every two different bases almost surely define different valuations (regardless of whether or not these valuations are solutions).
To this end, we can select two arbitrary different sets of $|V|$ edges that define potential bases $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$; they can be selected before drawing $\mathcal{G}^{\prime}$.
What we show is that it happens with probability 00 that $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ are bases and define the same solution in $\mathcal{G}$. If $\mathbf{b}_{1}$ or $\mathbf{b}_{2}$ is not a basis (which would happen if the inequations are not linearly independent, e.g. if a vertex $v$ does not appear at all in these inequations) we are done, too.
As $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ are different, there will be one inequation that corresponds to an edge $e=(v,v^{\prime})$ that occurs in $\mathbf{b}_{2}$, but not in $\mathbf{b}_{1}$. As all weight disturbances are drawn independently, we assume without loss of generality that the weight disturbance to this edge is drawn last.
Now, the valuation $\mathsf{val}_{1}$ defined by $\mathbf{b}_{1}$ does not depend on this final draw. For $\mathsf{val}_{1}$, there is a value $w_{e}^{\prime}=\mathsf{val}_{1}(v)-\lambda_{e}\mathsf{val}_{1}(v^{\prime})$ that defines the weight $w_{e}^{\prime}$ that $e$ would need to have, in order to make the inequation sharp for $e$.
For the valuation $\mathsf{val}_{1}$ defined by $\mathbf{b}_{1}$ to be equal to the valuation $\mathsf{val}_{2}$ defined by $\mathbf{b}_{2}$, the weight for the edge $e$ (after adding the drawn distortion) needs to be exactly $w_{e}^{\prime}$. (This is a necessary condition, but not necessarily a sufficient condition.) There is at most one value for the disturbance that would provide for this, and this disturbance for the weight of $e$ is sampled with a likelihood of 00.
Putting these two results together, we get the following:
@@
Given a pair of DPGs $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ with contraction $\lambda^{*}$ and gap $\gamma$, and $\mathcal{G}^{\prime}=(V_{\min},V_{\max},E,w^{\prime},\lambda)$ obtained from $\mathcal{G}$ by adding independently uniformly at random drawn values from an interval $(-\varepsilon,\varepsilon)$ to every edge weight, for some $\varepsilon<\frac{1-\lambda^{*}}{2}\gamma$, it holds that any joint co-optimal strategy from $\mathcal{G}^{\prime}$ is also co-optimal for $\mathcal{G}$, and $\mathcal{G}^{\prime}$ is almost surely sharp.
Note that we can estimate the gap cheaply when all coefficients in $\mathcal{G}$ are rational. To estimate the gap
$\gamma=\min\{-\mathsf{offset}(\mathsf{val}_{\sigma},e)\mid\sigma\mbox{ is a non-co-optimal joint strategy},e\in E,\mathsf{offset}(\mathsf{val}_{\sigma},e)<0\}\ ,$
we first fix a joint non-co-optimal strategy $\sigma$ and an edge $e=(v,w)$ with $\mathsf{offset}(\mathsf{val}_{\sigma},e)<0$.
Starting in $v$, $\sigma$ defines a run $\rho=e_{1}e_{2}e_{3}\ldots$ in the form of a “lasso path”, which consists of a (possibly empty) initial path $e_{1},\ldots,e_{k}$, followed by an infinitely often repeated cycle $e_{1}^{\prime},\ldots,e_{\ell}^{\prime}$, where the only vertex that occurs twice on the path $e_{1},\ldots,e_{k},e_{1}^{\prime},\ldots,e_{\ell}^{\prime}$ is the vertex reached before $e_{1}^{\prime}$ and after $e_{\ell}^{\prime}$, while all other vertices occur only once.
In this run, an edge $e_{i}$ occurs only once and contributes $(\prod_{j=1}^{i-1}\lambda_{e_{j}})w_{e_{i}}$ to the value of this run, while an edge $e_{i}^{\prime}$ occurs infinitely often and contributes the value $\frac{(\prod_{j=1}^{k}\lambda_{e_{j}})\cdot(\prod_{j=1}^{i-1}\lambda_{e_{j}^{\prime}})\cdot w_{e_{i}^{\prime}}}{1-\prod_{j=1}^{\ell}\lambda_{e_{j}^{\prime}}}$ to the value of the run. Now all we need to do is to estimate a common denominator.
To do this, let $\mathsf{nom}(r)$ and $\mathsf{denom}(r)$ be the nominator and denominator of a rational number $r$. It is easy to see that
@@
is a common denominator of all expressions of the first type, ($\prod_{j=1}^{i-1}\lambda_{e_{j}})w_{e_{i}}$, and of the nominator from the second, $(\prod_{j=1}^{k}\lambda_{e_{j}})\cdot(\prod_{j=1}^{i-1}\lambda_{e_{j}^{\prime}})\cdot w_{e_{i}^{\prime}}$.
This leaves the contribution of the denominator of the fraction from the second form, which is the nominator of the term $1-\prod_{j=1}^{\ell}\lambda_{e_{j}^{\prime}}$. But since the term is between 00 and $1$, its value is strictly smaller than the value of its denominator; we thus have
$\mathsf{nom}(1-\prod_{j=1}^{\ell}\lambda_{e_{j}^{\prime}})<\prod_{j=1}^{\ell}\mathsf{denom}(\lambda_{e_{j}^{\prime}})\ .$
To obtain $-\mathsf{offset}(\mathsf{val}_{\sigma},e)$, we have to compare with a run starting in $v$, but taking the edge $e\neq(v,\sigma(v))$ first.
This defines a run $\rho^{\prime}=d_{1}d_{2}d_{3}\ldots$ (with $d_{1}=e$), again in the form of a “lasso path” that consists of an initial path $d_{1},\ldots,d_{k^{\prime}}$, followed by an infinitely often repeated cycle $d_{1}^{\prime},\ldots,d_{\ell^{\prime}}^{\prime}$, where the only vertex that occurs twice on the path $d_{1},\ldots,d_{k},d_{1}^{\prime},\ldots,d_{\ell^{\prime}}^{\prime}$ is the vertex reached before $d_{1}^{\prime}$ and after $d_{\ell^{\prime}}^{\prime}$, while all other vertices occur only once.
In this run, an edge $d_{i}$ occurs only once and contributes $(\prod_{j=1}^{i-1}\lambda_{d_{j}})w_{d_{i}}$ to the value of this run, while an edge $d_{i}^{\prime}$ occurs infinitely often and contributes the value $\frac{(\prod_{j=1}^{k^{\prime}}\lambda_{d_{j}})\cdot(\prod_{j=1}^{i-1}\lambda_{d_{j}^{\prime}})\cdot w_{d_{i}^{\prime}}}{1-\prod_{j=1}^{{\ell^{\prime}}}\lambda_{d_{j}^{\prime}}}$ to the value of the run.
It is easy to see that
$\mathsf{common^{\prime}}\cdot\mathsf{denom}(\lambda_{e})\cdot\mathsf{denom}(w_{e})$
is a common denominator of all expressions of the first type, ($\prod_{j=1}^{i-1}\lambda_{d_{j}})w_{d_{i}}$, and of the nominator from the second, $(\prod_{j=1}^{k^{\prime}}\lambda_{d_{j}})\cdot(\prod_{j=1}^{i-1}\lambda_{d_{j}^{\prime}})\cdot w_{d_{i}^{\prime}}$.
This again leaves the contribution of the denominator of the fraction from the second form, which is the nominator of the term $1-\prod_{j=1}^{{\ell^{\prime}}}\lambda_{d_{j}^{\prime}}$. Again a term between 00 and $1$, its value is strictly smaller than the value of its denominator; we thus have
$\mathsf{nom}(1-\prod_{j=1}^{{\ell^{\prime}}}\lambda_{d_{j}^{\prime}})<\prod_{j=1}^{\ell^{\prime}}\mathsf{denom}(\lambda_{d_{j}^{\prime}})\ .$
The two loops are either the same or disjoined, and in both cases they have a common multiplier smaller than $\prod\limits_{v\in V}\mathsf{denom}(\lambda_{(v,\sigma(v))})$.
Obviously, the nominator will differ by at least one, so that $\gamma$ is at least the reciprocal of the largest such value.
This value is easily estimated by using the highest possible denominators available in $\mathcal{G}$, and its representation is polynomial in the size of $\mathcal{G}$.
@@
Let $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ be an improving DPG for a given set of positive numbers $\{\alpha_{e}>0\mid e\in E\}$, $\sigma$ a strategy for both players, $\mathsf{val}$ an optimal solution returned at Line 5 of Algorithm 3 for the adjusted function $f_{\sigma}^{\prime}$, and let there be no local improvements of $\sigma$ for $\mathsf{val}$. Then there is a neighbouring valuation $\mathsf{val}^{\prime\prime}$ to $\mathsf{val}$ such that there is a better strategy $\sigma^{\prime}$ that satisfies $f_{\sigma^{\prime}}(\mathsf{val}^{\prime\prime})<f_{\sigma}(\mathsf{val})$.
This strategy $\sigma^{\prime}$ can be selected in such a way that $(v,\sigma^{\prime}(v))\in E_{\mathsf{val}}^{\sigma}$ holds for all $v\in V$ for a given set $E_{\mathsf{val}}^{\sigma}$.
The proof of Theorem 5 can be repeated verbatim with the new offset definition.
@@
If each offset factor $\alpha_{e}$ is selected independently uniformly at random from a bounded interval of positive numbers777or from any other distribution over positive numbers that has 00 weight for all individual points, then a sharp DPG $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ is almost surely improving for a sampled set of positive numbers $\{\alpha_{e}>0\mid e\in E\}$.
@@
We show the claim that, for two arbitrary but fixed (before drawing the $\alpha_{e}$) different valuations $\mathsf{val}_{1}$ and $\mathsf{val}_{2}$ that satisfy the inequations in $H$ and any joint strategy $\sigma$, $f_{\sigma}(\mathsf{val}_{1})\neq f_{\sigma}(\mathsf{val}_{2})$ holds almost surely.
As $\mathcal{G}$ is sharp, this implies for all joint strategies $\sigma$ and all neighbouring bases $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ that define a solution and the valuations $\mathsf{val}_{1}$ and $\mathsf{val}_{2}$, respectively, that they define that $f_{\sigma}(\mathsf{val}_{1})\neq f_{\sigma}(\mathsf{val}_{2})$ holds almost surely. (Note that sharpness implies $\mathsf{val}_{1}\neq\mathsf{val}_{2}$ in this case.) That is, it implies that $\mathcal{G}$ is improving.
@@
We therefore have that $\mathsf{offset}(\mathsf{val}_{1},(v,\sigma(v)))\neq\mathsf{offset}(\mathsf{val}_{2},(v,\sigma(v)))$ holds for some $v\in V$. We fix such a vertex $v$.
As the $\alpha_{e}$ are drawn independently, we can assume w.l.o.g. that $\alpha_{(v,\sigma(v))}$ is drawn last. There is at most one value $\alpha_{(v,\sigma(v))}^{\prime}$ for which the following condition is not satisfied:
@@
It therefore holds almost surely that $f_{\sigma}(\mathsf{val}_{1})\neq f_{\sigma}(\mathsf{val}_{2})$.